OpenStudy (3psilon):

\[\ c-5 = \sqrt(4c+1)\]

OpenStudy (3psilon):

What is the value of C ? Help plz

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

\[(c-5)^2=4c+1\]

OpenStudy (anonymous):

\[c^2-10c+25=4c+1\]

OpenStudy (3psilon):

I got 2 and 12 but it just says it's 12

OpenStudy (anonymous):

\[c^2-14c+24=(c-12)(c-2)\]

OpenStudy (anonymous):

\[\sqrt{4(2)+1}\neq2-5\]

OpenStudy (3psilon):

? What does that mean?

OpenStudy (3psilon):

ohhh Can't have a negative root . Got it

OpenStudy (anonymous):